OpenStudy (anonymous):

If you were asked to convert 25 kg to the unit mg, which of the following would be the first fraction used in the conversion?

OpenStudy (anonymous):

\[25 kg = 25kg \times \frac{ 1000g }{ kg }\times \frac{ 1000mg }{ g}\]